Hi I asked this question before but forgot something How do you figure out the area of an triangle with all sides equal and the only number given is 5 thanks
Answer by tutorcecilia(2152)

Since all sides are equal and I am assuming that one side = 5, than all sides =5 including the base.
Draw a picture of the triangle with all sides =5.
.
Area=(1/2)(base)(height)[Formula for the area of a triangle]
A=(1/2)(5)(h) [Plug-in the values]
.
To find the value of the height, use the Pythagorean Theorem:
a^2+b^2=c^2
Since entire base is 5, use (1/2)(base)=(1/2)(5)=2.5 in the Theorem:
.
(2.5)^2+b^2=c^2
.
Since all sides=5, use 5 as one side in the Theorem:
(2.5)^+5^2=c^2
6.25+25=c^2
31.25=c^2
sqrt(31.25)=sqrt(C^2)
5.59=c
5.59=height
.
A=(1/2)bh [Plug-in the values]
A=(1/2)(5)(5.59)
A=13.9754
.
Check by plugging all of the values back into the Pythagorean Theorem.